OpenStudy (anonymous):

Consider the parametric equations x=cos t - sin t y=cos t + sin t 0

OpenStudy (amistre64):

well, it might be helpful to know that r^2 = x^2+y^2 other than that its hard to see pulling out a t at the moment

OpenStudy (amistre64):

what ideas have you tried so far?

OpenStudy (amistre64):

as far as plotting goes, its a circle ...

OpenStudy (sirm3d):

add and subtract the two equations: \[x+y=2\cos t\\y-x=2\sin t\]

OpenStudy (sirm3d):

square both equations, add them, and use a trigonometric identity.

OpenStudy (amistre64):

x = cos t - sin t ; y = cos t + sin t x^2 = cos^2 t -2sin t cos t + sin^2 t y^2 = cos^2 t +2sin t cos t + sin^2 t ------------------------------- r^2 = 2(cos^2 t + sin^2 t) = 2 x^2 + y^2 = 2

OpenStudy (amistre64):

yeah, thats what i kept coming up with :) didnt know if the question needed a function tho